

********************************************************************
UARS WEEKLY STATUS REPORT - 18 OCT 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
DR. DIANNE PRINZ
On Saturday, October 11, Dr. Dianne Prinz, SUSIM Principal
Investigator from 1998 to 2000, passed away at her home after a
long illness.
Dianne received her B.S. degree from the University of Pittsburgh
in 1960 and her Ph.D. from the Johns Hopkins University in 1967.
Among her many accomplishments was the development of a Lyman
alpha spectroheliograph which provided high resolution images of
the Sun. Later, she trained to become a astronaut (payload
specialist), a career path which was cut short by the Space
Shuttle Challenger accident in 1986.
During SUSIM's development phases, Dianne provided the optical
design for the instrument. Dianne always emphasized the need for
the thorough understanding of the instrument's functioning. Her
guidance during the SUSIM experiment's operational phase was
crucial in making SUSIM the success that it is today.
All of us on the SUSIM team have fond memories of Dianne and will
miss her very much.
SUSIM was switched off as a part of the yaw-around for the period
Oct. 10-14. Otherwise, SUSIM successfully performed its daily
mid-resolution and MgII solar scans during periods of predicted
UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 September 2002 (Orbit 60520, GMT Day 278)
through 11 October 2002 (Orbit 60623, GMT Day 284).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 279/2202 60548 POWER OFF AXIS
HALOE 280/2208 60563 POWER OFF
SUSIM 282/2159 60593 POWER OFF
PEM 282/2159 60593 POWER OFF ZEPS/HEPS
SOLSTICE 282/2159 60593 DAY ONLY MODE
HRDI 283/0220 60595 POWER OFF
A Forward-to-Reverse Yaw Around maneuver (number 113) was successfully
executed in Orbit 60604 beginning 283/15:49:12Z and completing at
283/16:26:22Z on 10 October 2002. All instruments were in their off
or standby mode (including HRDI and PEM). Due to the SA position at
269 degrees, the effective night length of the spacecraft after the
maneuver was approximately 56.80 minutes. Post maneuver analysis of
data indicate that during the Yaw Around, the end-of-night load bus
voltage was 26.08 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 20.74 and 20.16 percent
respectively. The next (reverse-to-forward) YAW is scheduled for 19
November 2002 (DOY 323).
The beta angle ranged between the angles of 23.3 degrees to 04.5
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 71 degrees on 31 October 2002 (DOY 304).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.64 to 0.96 degrees C. Additional
battery performance data for each battery on 04 October and 11 October
2002 are provided in Appendix D.
The clock error ranged from -10.1 to 10.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 03
PEM = 02
HALOE = 01
SUSIM = 01
SOLSTICE = 01
HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
Two (2) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The new hard disk on the prototype was installed as well as the new
build of the Eclipse software. Once the testing of this new build is
complete the software will be released for operations. This will most
likely not take place until the ERBS integration is complete.
The FOT is also working hard in preparation for the college students
that will be in the control center to learn about spacecraft
operations. The FOT will try to formulate training materials and
structure a learning plan to teach the students. This is all part of
the new Educational Outreach initiative/contract, which started on
October 1, 2002 for both UARS and TOMS.
MEETINGS
- The FOT attended a meeting on 09 October 2002 to discuss the
FOT's plan for training the students.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 19 November 2002 (DOY 323).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT
Occurrences: Appendix C, Clock & Solar Array Rate
Corrections: Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 October 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
60548 279/2202 02-345 PEM POWER OFF AXIS
60563 280/2208 02-347 HALOE POWER OFF
60593 282/2159 02-348 SUSIM POWER OFF
60593 282/2159 02-349 PEM POWER OFF ZEPS/HEPS
60593 282/2159 02-350 SOLSTICE DAY ONLY MODE
60595 283/0220 02-351 HRDI POWER OFF (INST PB)
60603 283/1420 02-352 MACS YAW MANEUVER
60607 283/2131 02-353 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
60618 284/1514 02-352 PSU VT LEVEL 6 TO VT LEVEL 5
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 October 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 October 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
60607 283/2134 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 October 2002
04 October 2002 - GMT Day 277 Beta = 27.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -140.0/-168.0 -3.05/-4.31 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-89.6 +5.15/+4.19 +28.8/-15.2 24.8 1.030 21.0
3 +22.4/-67.2 +4.19/+3.23 +28.0/-15.2 24.8 1.020 20.5
11 October 2002 - GMT Day 284 Beta = 04.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -145.6/-168.0 -3.68/-4.63 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-33.6 +5.46/+4.50 +28.8/-13.6 26.2 1.020 20.3
3 +22.4/-33.6 +3.87/+3.23 +27.2/-13.2 26.1 1.020 19.6
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
60603 283/1430 L5 -> L6
60618 284/1538 L6 -> L5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 October 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-051 FOT 60366 267/1951 30832 GENERIC LATE ACQUISITION
02-052 SOLSTICE 60366 267/1951 N/A LOADING PROBLEM